I'm stuck on one part of deriving the quadratic equation. I can not figure out how x^2 + (b/a)x + (b/2a)^2 becomes (x + b/2a)^2. If I multiply this out I get x^2 + (b/2a)x + (b/2a)x + (b/2a)^2.
Thanks

Answer by mukku(2)

(a+b)^2 = a^2+b^2+2ab
Here a ix and b = b/2a
Substituting in the formula foe (a+b)^2 we get what you have in your problem